OpenStudy (anonymous):

use the generalized power rule to find the derivative of each function. f(x)=2x(x3-1)4 The 3 and 4 are powers

OpenStudy (anonymous):

\[\large{f' (x)=2(x^3-1)^4 + 8x(x^3-1)^3 \times 3x^2 }\]\[=\large{2(x^3-1)^4 + 24x^3(x^3-1)^3}\]